OpenStudy (anonymous):

Find the general solution of the differential equation. Use C for the constant of integration. dr/dp = 10sin(p)

OpenStudy (anonymous):

would it be -10xcosp+C..?

OpenStudy (darkprince14):

integrate poth sides wrt p.

OpenStudy (unklerhaukus):

that's it

OpenStudy (anonymous):

okay let me try it! hang on!!

OpenStudy (anonymous):

it was wrong :/

OpenStudy (darkprince14):

\[dr = 10 \sin (p) dp\]\[\int\limits dr = \int\limits 10 \sin(p)dp\] then work your magic :)

OpenStudy (anonymous):

i still get -10cosp + C.. how would you go about it..? am i forgetting something

OpenStudy (darkprince14):

well, actually that's correct. Take a look at the first answer you've posted and you'll see the difference :)

OpenStudy (anonymous):

i forgot to capitalize the C that why i was getting it wrong -_- thank you! lol

OpenStudy (darkprince14):

you got x after 10, you wrote -10xcosp...

OpenStudy (anonymous):

yeah that too. Thanks! :)

OpenStudy (unklerhaukus):

i thought you meant `x` for times

OpenStudy (darkprince14):

or is that the multiplication sign? lol

OpenStudy (darkprince14):

just realized that XD